# FUSAR-KLIP: Towards Multimodal Foundation Models for Remote Sensing

### v2

@@ 3.3 FUSAR-KLIP Overall Framework
Specifically, FUSAR-KLIP adopts a standardized dual-encoder architecture to separately model visual and textual representations. The visual encoder employs a vision transformer, where SAR images are divided into fixed-size patches and processed through linear projection and positional encoding before being passed to the Transformer. This yields a uniform visual embedding vector: $f_{v}\in\mathbb{R}^{d}$. The text encoder is based on BERT, which tokenizes and embeds the input text, producing a corresponding embedding vector: $f_{t}\in\mathbb{R}^{d}$ through a multi-layer Transformer. These embeddings are aligned within a shared semantic space. ViT and BERT, both Transformer-based, are widely adopted in multimodal frameworks (e.g., BLIP, ALBEF, OFA[55],[56]) due to their strong alignment capability, generalization, and downstream task compatibility, making them well-suited as foundational encoders for our model.
@@ 8 Phase 2: Text Filtering and Refinement
$\small\Delta\mathcal{L}_{\text{ITC}}(t_{i})=\mathcal{L}_{\text{ITC}}(T,I)-\mathcal{L}_{\text{ITC}}(T\setminus t_{i},I)$
if $\Delta\mathcal{L}_{\text{ITC}}(t_{i})>0$ then
@@ 4.1 Implementation Details
ChangeClip[58] ViT-Base ISPRS-2024 59,246
@@
Geochat[45] ViT-Large CVPR-2024 141,246
VHM[43] ViT-Large AAAI-2025 1,390,405
@@
CLIP*[60] ViT-Base ICML-2021 400 M
BLIP*[61] ViT-Base ICML-2022 100 M
@@ 4.2 Analysis of FUSAR-GEOVL-1M Text
During dataset construction, we designed task-oriented prompts based on the Hierarchical Cognitive Chain-of-Thought (HCoT) framework introduced in Section 3.2, aiming to enhance the reasoning capability and fine-grained perception of large language models in SAR image interpretation. Instruction sets were developed for three representative RS tasks: ground target recognition, marine target recognition, and terrain understanding. For target-related tasks, prompts guide the model to analyze attributes such as location, quantity, category, and structural features. In terrain-oriented tasks, the focus is on interpreting elements such as buildings, water bodies, and road networks. Each SAR image is paired with eight complementary textual descriptions spanning multiple semantic levels and knowledge dimensions. Key information is extracted via regular expression matching, resulting in a total of one million structured descriptions.
@@
To assess the textual accuracy of FUSAR-GEOVL-1M, we conducted an expert evaluation on a randomly sampled 2% subset of the dataset, given its large scale. Each image was annotated with eight types of semantic information, and accuracy was independently assessed for each type. As shown in Fig. 9, the overall accuracy approaches 80%. Marine target descriptions achieve higher accuracy than ground targets, benefiting from the relatively uniform sea background. Environmental descriptions also perform well, aided by geospatial cues and lower semantic complexity. While fine-grained target attributes occasionally contain errors, the generated texts are generally accurate and coherent, offering strong support for high-level scene understanding and multimodal pre-training.
@@
To further compare semantic expressiveness, we selected a subset of FUSAR-GEOVL-1M with manual target annotations and generated templated descriptions using the method in RemoteCLIP [13]. As shown in Fig. 10, t-SNE visualization reveals that FUSAR-GEOVL-1M texts exhibit tighter clustering aligned with image features, whereas templated texts are more dispersed, indicating semantic sparsity. Additionally, as illustrated in Fig. 11, we compared and analyzed multiple RS multimodal datasets using two metrics: the text semantic richness index (MTLD[62]) and token count. Overall, FUSAR-GEOVL-1M significantly outperforms existing mainstream datasets in both semantic richness and token count.
@@ 4.3 Visual Task Benchmarks
All models use the visual encoder weights in the multimodal foundation model as initialization parameters for subsequent SFT. In addition, three single-modal VIT foundation models are also used for experimental comparison. Among them, MAE[66] and MOCO v3[67] are pre-trained based on the dataset proposed in this paper, while SAR-JEPA[37] is the foundation model proposed by Li et al. for SAR image interpretation. Table III summarizes the key model configurations and training hyperparameters of each subtask. The training and test data are divided in a ratio of 4:1 by default.
@@
Classification ViT-Cls[31] AdamW 0.003 100 Warmup,CosineDecay
Detection ViTDET[68] AdamW 0.0001 36 Warmup,MultiStep
Segmentation Segmenter[69] AdamW 0.001 100 PolyLR
@@
1) Target Classification: The classifier is implemented using the PyTorch framework and evaluated on three representative SAR datasets: FUSAR-SHIP [22], FUSAR-AIR [54], and SAR-ACD [21]. FUSAR-SHIP and FUSAR-AIR are curated subsets of FUSAR-GEOVL-1M, focusing on ship and aircraft target recognition, respectively. FUSAR-SHIP, developed by the Key Laboratory of Electromagnetic Wave Information Science at Fudan University, includes 5,242 images spanning 15 ship categories and 98 subcategories. FUSAR-AIR comprises diverse aircraft types such as transport, refueling, and civil aircraft. SAR-ACD, a public third-party dataset with higher scene complexity and 1-meter resolution, contains 3,032 images across six aircraft types, with approximately 500 samples per class.
@@
2) Target Detection: This study developed an object detection model based on the typical ViT detection framework, ViTDet. By combining the ViT backbone with the feature pyramid structure, ViTDet enhances the detection of multi-scale targets while maintaining global modeling capabilities, achieving excellent performance across various visual detection tasks. We evaluated the model’s performance using three representative SAR image datasets, under the AP@50 metric. FUSAR-SHIP-Sense (FU-SS) consists of FUSAR-SHIP target slices expanded into complete scene images, totaling 3,838 images. FUSAR-AIR-Sense (FU-AS) is composed of FUSAR-AIR target slices expanded into complete scene images, totaling 2,491 images. SAR-AIRcraft-Few (AIR-F)[26] is an open-source aircraft detection dataset built from Gaofen-3 images, containing 4,368 images, 16,463 targets, and 7 aircraft categories. This experiment used only 20% of the training data to assess the generalization ability of the pre-trained model in non-homologous and few-sample scenarios.
@@
3) Segmentation: This study constructed a SAR image semantic segmentation model based on PyTorch, employing ViT as the visual encoder and the Mask Transformer from Segmenter as the decoder to enhance semantic modeling. The model was evaluated using overall accuracy (OA) and mean Intersection over Union (mIoU) on two representative datasets. FUSAR-MAP[16] focuses on fine-grained segmentation of urban scenes—including buildings, roads, vegetation, and background—with 600 high-resolution SAR images (1024×1024). AIR-PolarSAR-Seg (PoSAR-Seg)[27] targets six land cover categories across 2,000 image slices (512×512), enabling evaluation of general semantic segmentation performance.
The experimental results are presented in Table IV. FUSAR-KLIP achieved the highest $OA$ and $mIoU$ on both segmentation datasets, significantly outperforming existing multimodal and unimodal methods. Among the multimodal models, performance varied greatly due to differences in modeling capabilities of RS domain knowledge. Models such as RemoteClip, GeoRSClip, CLIP, and SkyScript performed relatively well. The unimodal model SAR-JEPA performed well on FUSAR-MAP but showed a significant decline in performance on the more complex AIR-PolarSAR-Seg.
@@ 4.4 Visual-Language Tasks Benchmarks
The experimental configuration is based on the fine-tuning settings of mainstream multimodal models, such as BLIP[61], used in the natural image domain, with adjustments made to suit the characteristics of the data. Table V summarizes the core training parameters for the three visual language tasks. Considering task-specific differences, the captioning task is configured with a larger maximum token length, while the VQA task uses a slightly higher learning rate to accelerate convergence.
@@
1) Image-Text Retrieval: This task evaluates the model’s cross-modal retrieval capability between SAR images and text in both directions: Image→Text and Text→Image, which is crucial for RS image retrieval. Standard $Recall@K$ metrics ($R@1$, $R@5$, $R@10$) are used to assess performance, with $txt\_{r1/r5/r10}$ and img_r1/r5/r10img\_r1/r5/r10 denoting the Top-K recall rates for text-to-image and image-to-text retrieval, respectively. The average of these six values serves as the overall retrieval metric. Quantitative and qualitative results are shown in Table VI and Fig. 14. As shown in Fig. 14, we compare the proposed method with a suboptimal model. FUSAR-KLIP has the highest matching degree for the correct option and can therefore retrieve the corresponding image more accurately.
2) Image Captioning: The Caption task aims to generate accurate descriptions of key targets and their spatial distribution in SAR images using natural language. We evaluate the quality of the image descriptions using four mainstream metrics[61]: BLEU-4, METEOR, CIDEr, and SPICE. These metrics assess the description quality from four dimensions: phrase matching, semantic consistency, content relevance, and structural integrity. Table VI presents the quantitative evaluation results for each pre-trained model on the Caption task. FUSAR-KLIP achieves leading performance across many indicators, particularly demonstrating significant advantages in CIDEr.
@@ 4.5 Comparison with general multimodal large models
Some commercial multimodal large models with ultra-large parameter volumes (such as GPT-4o, Gemini 2.5, and Grok-3) have demonstrated strong multi-task capabilities. However, due to their closed-source nature, their applicability in the RS field still requires further verification. Given that VQA is the most common application form for multimodal language models, we use the VQA task to assess the reasoning ability of these models on SAR images. The experimental results are shown in Table VII. The general large models perform the weakest in the target counting task, and their accuracy in other tasks is significantly lower compared to the model proposed in this study. This indicates that general multimodal large models suffer from insufficient adaptability in SAR tasks, while the foundational model developed in this research provides robust support for advancing multimodal models in the SAR domain.
@@
GPT-4o[30] 33 66 28 53 47 44
Gemini-2.5[70] 12 38 43 56 73 38
Grok-3[71] 19 27 54 33 65 48
Qwen2.5-VL[72] 36 47 29 51 33 39
@@ 4.6 Ablation experiment
In the data validation, we selected the latest large-scale SAR text dataset, SARLANG[14], for comparison. This dataset depends on target detection annotations and fixed templates to generate image-text pairs, ensuring high semantic accuracy without the need for additional screening or reconstruction. However, as analyzed in Fig. 11, SARLANG exhibits issues such as sparse semantics and limited knowledge coverage.
@@ 5 Conclusion
In this work, we present FUSAR-KLIP, the first general-purpose multimodal foundation model designed for SAR. To support its development, we construct FUSAR-GEOVL-1M, the largest SAR image–text dataset to date and the first to comprehensively preserve geographic attributes. Distinct from prior datasets, FUSAR-GEOVL-1M achieves spatial scale consistency, cross-platform adaptability, and hierarchical semantic representation. By further integrating HCoT and SCIO, our framework offers a systematic solution to enhance cross-modal alignment and knowledge integration.
We establish a unified SAR multimodal benchmark covering 11 representative downstream tasks across both vision and language domains, and compare FUSAR-KLIP against 14 leading foundation models. The results consistently demonstrate FUSAR-KLIP’s superior performance, particularly in challenging tasks such as object counting and land-cover classification, highlighting its ability to advance beyond perceptual-level recognition toward semantic-level cognition.
This study systematically demonstrates the feasibility of extending self-supervised learning to SAR multimodal modeling and introduces a paradigm that incorporates geographic priors and cognitive reasoning. Looking forward, we plan to further enlarge the dataset scale, enrich task diversity, and explore unified modeling across multiple modalities and domains. Ultimately, this line of research aims to build a comprehensive multimodal foundation model system for remote sensing, empowering large-scale Earth observation and intelligent interpretation.

### v3

@@ 3.3 FUSAR-KLIP Overall Framework
Specifically, FUSAR-KLIP adopts a standardized dual-encoder architecture to separately model visual and textual representations. The visual encoder employs a vision transformer, where SAR images are divided into fixed-size patches and processed through linear projection and positional encoding before being passed to the Transformer. This yields a uniform visual embedding vector: $f_{v}\in\mathbb{R}^{d}$. The text encoder is based on BERT, which tokenizes and embeds the input text, producing a corresponding embedding vector: $f_{t}\in\mathbb{R}^{d}$ through a multi-layer Transformer. These embeddings are aligned within a shared semantic space. ViT and BERT, both Transformer-based, are widely adopted in multimodal frameworks (e.g., BLIP, ALBEF, OFA[li2021align],[Liu2025Causal]) due to their strong alignment capability, generalization, and downstream task compatibility, making them well-suited as foundational encoders for our model.
@@ 8 Phase 2: Text Filtering and Refinement
$\small\Delta\mathcal{L}_{\text{ITC}}(t_{i})=\mathcal{L}_{\text{ITC}}(T-t_{i},I)-\mathcal{L}_{\text{ITC}}(T,I)$
if $\Delta\mathcal{L}_{\text{ITC}}(t_{i})<0$ then
@@ 4.1 Implementation Details
ChangeClip[dong2024changeclip] ViT-Base ISPRS-2024 59,246
@@
Geochat[kuckreja2024geochat] ViT-Large CVPR-2024 141,246
VHM[pang2025vhm] ViT-Large AAAI-2025 1,390,405
@@
[…]
CLIP*[radford2021learning] ViT-Base ICML-2021 400 M
BLIP*[li2022blip] ViT-Base ICML-2022 100 M
@@ 4.2 Analysis of FUSAR-GEOVL-1M Text
During dataset construction, we designed task-oriented prompts based on the HCoT framework introduced in Section 3.2, aiming to enhance the reasoning capability and fine-grained perception of large language models in SAR image interpretation. Instruction sets were developed for three representative remote sensing tasks: ground target recognition, marine target recognition, and terrain understanding. For target-related tasks, prompts guide the model to analyze attributes such as location, quantity, category, and structural features. In terrain-oriented tasks, the focus is on interpreting elements such as buildings, water bodies, and road networks. Each SAR image is paired with eight complementary textual descriptions spanning multiple semantic levels and knowledge dimensions. Key information is extracted via regular expression matching, resulting in a total of one million structured descriptions.
@@
To assess the textual accuracy of FUSAR-GEOVL-1M, we conducted an expert evaluation on a randomly sampled 2% subset of the dataset. Three experts independently rated eight semantic attributes per image, assigning information accuracy scores to quantify factual consistency. As shown in Fig. 9, the overall accuracy approaches 80%. Marine target descriptions achieve higher accuracy than ground targets, benefiting from the relatively uniform sea background. Environmental descriptions also perform well, aided by geospatial cues and lower semantic complexity. While fine-grained target attributes occasionally contain errors, the generated texts are generally accurate and coherent, offering strong support for high-level scene understanding and multimodal pre-training.
@@
To further compare semantic expressiveness, we selected a subset of FUSAR-GEOVL-1M with manual target annotations and generated templated descriptions using the method in RemoteCLIP [10504785]. As shown in Fig. 10, t-SNE visualization reveals that FUSAR-GEOVL-1M texts exhibit tighter clustering aligned with image features, whereas templated texts are more dispersed, indicating semantic sparsity. Additionally, as illustrated in Fig. 11, we compared and analyzed multiple remote sensing multimodal datasets using two metrics: the text semantic richness index (MTLD[mccarthy2005assessment]) and token count. Overall, FUSAR-GEOVL-1M significantly outperforms existing mainstream datasets in both semantic richness and token count.
@@ 4.3 Visual Task Benchmarks
All models use the visual encoder weights in the multimodal foundation model as initialization parameters for subsequent SFT. In addition, three single-modal VIT foundation models are also used for experimental comparison. Among them, MAE[he2022masked] and MOCO v3[Chen_2021_ICCV] are pre-trained based on the dataset proposed in this paper, while SAR-JEPA[LI2024326] is the foundation model proposed by Li et al. for SAR image interpretation. Table III summarizes the key model configurations and training hyperparameters of each subtask. The training and test data are divided in a ratio of 4:1 by default.
@@
Classification ViT-Cls[dosovitskiy2021an] AdamW 0.003 100 Warmup,CosineDecay
Detection ViTDET[li2022exploring] AdamW 0.0001 36 Warmup,MultiStep
Segmentation Segmenter[strudel2021segmenter] AdamW 0.001 100 PolyLR
@@
1) Target Classification: The classifier is implemented using the PyTorch framework and evaluated on three representative SAR datasets: FUSAR-SHIP [hou2020fusar], FUSAR-AIR [Qian2024Thesis], and SAR-ACD [9754573]. FUSAR-SHIP and FUSAR-AIR are curated subsets of FUSAR-GEOVL-1M, focusing on ship and aircraft target recognition, respectively. FUSAR-SHIP, developed by the Key Laboratory of Electromagnetic Wave Information Science at Fudan University, includes 5,242 images spanning 15 ship categories and 98 subcategories. FUSAR-AIR comprises diverse aircraft types such as transport, refueling, and civil aircraft. SAR-ACD, a public third-party dataset with higher scene complexity and 1-meter resolution, contains 3,032 images across six aircraft types, with approximately 500 samples per class.
@@
2) Target Detection: This study developed an object detection model based on the typical ViT detection framework, ViTDet. By combining the ViT backbone with the feature pyramid structure, ViTDet enhances the detection of multi-scale targets while maintaining global modeling capabilities, achieving excellent performance across various visual detection tasks. We evaluated the model’s performance using three representative SAR image datasets, under the AP@50 metric. FUSAR-SHIP-Sense (FU-SS) consists of FUSAR-SHIP target slices expanded into complete scene images, totaling 3,838 images. FUSAR-AIR-Sense (FU-AS) is composed of FUSAR-AIR target slices expanded into complete scene images, totaling 2,491 images. SAR-AIRcraft-Few (AIR-F)[Wang2023SARAIRcraft] is an open-source aircraft detection dataset built from Gaofen-3 images, containing 4,368 images, 16,463 targets, and 7 aircraft categories. This experiment used only 20% of the training data to assess the generalization ability of the pre-trained model in non-homologous and few-sample scenarios.
@@
3) Segmentation: This study constructed a SAR image semantic segmentation model based on PyTorch, employing ViT as the visual encoder and the Mask Transformer from Segmenter as the decoder to enhance semantic modeling. The model was evaluated using overall accuracy (OA) and mean Intersection over Union (mIoU) on two representative datasets. FUSAR-MAP[shi2021object] focuses on fine-grained segmentation of urban scenes—including buildings, roads, vegetation, and background—with 600 high-resolution SAR images (1024×1024). AIR-PolarSAR-Seg (PoSAR-Seg)[wang2022air] targets six land cover categories across 2,000 image slices (512×512), enabling evaluation of general semantic segmentation performance.
The experimental results are presented in Table IV. FUSAR-KLIP achieved the highest $OA$ and $mIoU$ on both segmentation datasets, significantly outperforming existing multimodal and unimodal methods. Among the multimodal models, performance varied greatly due to differences in modeling capabilities of remote sensing domain knowledge. Models such as RemoteClip, GeoRSClip, CLIP, and SkyScript performed relatively well. The unimodal model SAR-JEPA performed well on FUSAR-MAP but showed a significant decline in performance on the more complex AIR-PolarSAR-Seg.
@@ 4.4 Visual-Language Tasks Benchmarks
The experimental configuration is based on the fine-tuning settings of mainstream multimodal models, such as BLIP[li2022blip], used in the natural image domain, with adjustments made to suit the characteristics of the data. Table V summarizes the core training parameters for the three visual language tasks. Considering task-specific differences, the captioning task is configured with a larger maximum token length, while the VQA task uses a slightly higher learning rate to accelerate convergence.
@@
1) Image-Text Retrieval: This task evaluates the model’s cross-modal retrieval capability between SAR images and text in both directions: Image→Text and Text→Image, which is crucial for remote sensing image retrieval. Standard $Recall@K$ metrics ($R@1$, $R@5$, $R@10$) are used to assess performance, with $txt\_{r1/r5/r10}$ and img_r1/r5/r10img\_r1/r5/r10 denoting the Top-K recall rates for text-to-image and image-to-text retrieval, respectively. The average of these six values serves as the overall retrieval metric. Quantitative and qualitative results are shown in Table VI and Fig. 14. As shown in Fig. 14, we compare the proposed method with a suboptimal model. FUSAR-KLIP has the highest matching degree for the correct option and can therefore retrieve the corresponding image more accurately.
2) Image Captioning: The Caption task aims to generate accurate descriptions of key targets and their spatial distribution in SAR images using natural language. We evaluate the quality of the image descriptions using four mainstream metrics[li2022blip]: BLEU-4, METEOR, CIDEr, and SPICE. These metrics assess the description quality from four dimensions: phrase matching, semantic consistency, content relevance, and structural integrity. Table VI presents the quantitative evaluation results for each pre-trained model on the Caption task. FUSAR-KLIP achieves leading performance across many indicators, particularly demonstrating significant advantages in CIDEr.
@@ 4.5 Comparison with general multimodal large models
Some commercial multimodal large models with ultra-large parameter volumes (such as GPT-5, Gemini 3, and Grok-4) have demonstrated strong multi-task capabilities. However, due to their closed-source nature, their applicability in the remote sensing field still requires further verification. Given that VQA is the most common application form for multimodal language models, we use the VQA task to assess the reasoning ability of these models on SAR images. The experimental results are shown in Table VII. The general large models perform the weakest in the target counting task, and their accuracy in other tasks is significantly lower compared to the model proposed in this study. This indicates that general multimodal large models suffer from insufficient adaptability in SAR tasks, while the foundational model developed in this research provides robust support for advancing multimodal models in the SAR domain.
@@
GPT-5[OpenAIChatGPT] 35 67 28 52 49 47
Gemini-3[google_gemini_web] 21 40 41 56 71 42
Grok-4[xai_grok4] 19 29 54 36 67 48
Qwen3-VL[bai2025qwen3vltechnicalreport] 38 51 27 54 34 40
@@ 4.6 Ablation experiment
In the data validation, we selected the latest large-scale SAR text dataset, SARLANG[wei2025sarlang], for comparison. This dataset depends on target detection annotations and fixed templates to generate image-text pairs, ensuring high semantic accuracy without the need for additional screening or reconstruction. However, as analyzed in Fig. 11, SARLANG exhibits issues such as sparse semantics and limited knowledge coverage.
@@ 5 Conclusion
This paper addresses the fundamental cognitive disconnect between general visual representations and remote sensing SAR interpretation logic in cross-modal artificial intelligence, proposing FUSAR-KLIP—the first knowledge-guided general multimodal SAR foundational model. To this end, we constructed FUSAR-GEOVL-1M, the largest and first large-scale SAR image dataset to date that fully integrates geographic projection attributes. Compared to traditional datasets, FUSAR-GEOVL-1M exhibits significant advantages in spatial scale consistency and cross-platform adaptability. Furthermore, through our proposed Hierarchical Cognitive Chain and Self-Consistent Iterative Optimization mechanism, it successfully transforms implicit geographic environment and scattering features into explicit structured knowledge, providing systematic physical and geographic semantic support for cross-modal alignment.
We established a unified evaluation benchmark across 11 typical downstream tasks and conducted a systematic comparison with 15 mainstream multimodal models. Experimental results show that FUSAR-KLIP demonstrates leading performance across multiple tasks, particularly achieving breakthrough improvements in target counting and land cover classification tasks that heavily rely on spatial reasoning. This result strongly validates that incorporating geographical priors and physical cognition into representation learning is a key path to overcoming the bottlenecks in remote sensing interpretation and bridging the gap between machine representation and human cognition.
This study establishes for the first time the core role of ”knowledge-guided” in the construction of remote sensing foundation models and proposes a new paradigm that integrates geographic priors and cognitive modeling. It also provides theoretical reference for modeling other remote sensing modalities with significant physical characteristics, such as multispectral and infrared. In the future, we will focus on expanding data scale and task types, further exploring unified modeling across modalities and domains, and moving towards building a more complete and logically consistent intelligent system for Earth observation.